the length of a rectangle is 6 more yards than the width. If the perimiter is 184 yards, what are the length and width

Answer by ewatrrr(24785)

Hi,
*Note: Let x be the width. Then (x+6) is the length.
In general, the Perimeter is 2w + 2l
.
2x + 2(x+6) = 184yds
.
simplify and solve
4x + 12 = 184yds
4x = 172yds
.
x = 43yds (the width)
length is 49yds
.
check your answer
86 + 2*49 = 184